# Dynamic Benchmark Construction for Evaluating Large Language Models on Real-World Codes

### v1

@@ H.3 WSC Python Example: calculate_ngram_repetition
Calculates the proportion of repeated n-grams in a given text.
[…]
text (str): The input text to analyze, consisting of words separated by spaces.
n (int): The size of the n-grams to generate (e.g., 2 for bigrams, 3 for trigrams).
[…]
float: The proportion of n-grams that are repeated in the text. Returns 0 if no n-grams can be generated.
[…]
ValueError: If n is less than or equal to 0.
# TODO: Implement this function
pass
Testcase Generator
from hypothesis import settings, given, Verbosity, example
from hypothesis import strategies as st
import json
import os
import atexit
import copy
from collections import Counter
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
os.makedirs(TEST_CASE_DIR, exist_ok=True)
TEST_CASE_FILE = os.path.join(TEST_CASE_DIR, "test_cases.json")
generated_cases = []
stop_collecting = False # Global flag to control case collection
[…]
def calculate_ngram_repetition(text, n):
words = text.split()
ngrams = [tuple(words[i : i + n]) for i in range(len(words) - n + 1)]
ngram_counts = Counter(ngrams)
total_ngrams = len(ngrams)
repeated_ngrams = sum(1 for count in ngram_counts.values() if count > 1)
return repeated_ngrams / total_ngrams if total_ngrams > 0 else 0
[…]
def text_strategy():
return st.text(
alphabet=st.characters(whitelist_categories=(’L’, N’, Z’), min_codepoint=32, max_codepoint=126),
min_size=0, max_size=100
[…]
return st.integers(min_value=1, max_value=5)
[…]
@settings(max_examples=10000, verbosity=Verbosity.verbose, print_blob=True)
@given(text=text_strategy(), n=n_strategy())
@example(text="", n=1)
@example(text="a", n=1)
@example(text="a b c", n=2)
@example(text="a a b b c c", n=2)
@example(text="a b c d e f", n=3)
@example(text="a a a a a a", n=3)
def test_calculate_ngram_repetition(text, n):
global stop_collecting
if stop_collecting:
return
[…]
text_copy = copy.deepcopy(text)
n_copy = copy.deepcopy(n)
[…]
try:
expected = calculate_ngram_repetition(text_copy, n_copy)
except Exception:
return # Skip inputs that cause exceptions
[…]
generated_cases.append({
"Inputs": {
"text": text_copy,
"n": n_copy
[…]
if len(generated_cases) >= 500:
stop_collecting = True
[…]
def save_test_cases():
with open(TEST_CASE_FILE, "w") as f:
json.dump(generated_cases, f, indent=2, ensure_ascii=False)
print(f"✅ Saved {len(generated_cases)} test cases to {TEST_CASE_FILE}")
[…]
Testcase Runner
import json
import os
import copy
from collections import Counter
from helper import deep_compare
from tested import calculate_ngram_repetition as func1
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
TEST_CASE_JSON_PATH = os.path.join(TEST_CASE_DIR, "test_cases.json")
[…]
def calculate_ngram_repetition(text, n):
words = text.split()
ngrams = [tuple(words[i : i + n]) for i in range(len(words) - n + 1)]
ngram_counts = Counter(ngrams)
total_ngrams = len(ngrams)
repeated_ngrams = sum(1 for count in ngram_counts.values() if count > 1)
return repeated_ngrams / total_ngrams if total_ngrams > 0 else 0
[…]
if not os.path.exists(TEST_CASE_JSON_PATH):
print(f"JSON file not found: {TEST_CASE_JSON_PATH}")
return []
with open(TEST_CASE_JSON_PATH, "r") as f:
test_cases = json.load(f)
return test_cases
[…]
# Use deep_compare for basic types (int, float, str, etc.)
return deep_compare(expected, actual, tolerance=1e-6)
[…]
for i, case in enumerate(test_cases):
inputs = copy.deepcopy(case["Inputs"])
text = inputs["text"]
n = inputs["n"]
[…]
expected_output = calculate_ngram_repetition(text, n)
actual_output = func1(text, n)
[…]
if compare_outputs(expected_output, actual_output):
print(f"Test case {i + 1} passed.")
else:
print(f"Test case {i + 1} failed:")
print(f" Inputs: {inputs}")
print(f" Expected: {expected_output}")
print(f" Actual: {actual_output}")
[…]
test_cases = load_test_cases_from_json()
run_tests_with_loaded_cases(test_cases)
H.4 Benchmark Runner Prompt
Table 16: Prompt Template for Python Benchmark Runner in Code2Bench
@@ H.4 Benchmark Runner Prompt
Table 17: Prompt Template for Weakly Self-Contained Benchmark Runner in Code2Bench
@@
Table 18: Prompt Template for Go Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write Go code.
## Input
The user provides you with an function signature and docstring, you should generate a Go function based on them.
## Output “‘go The generated Go code. “‘
## Note
Provide only Go code within a “‘go“‘ code block, including a complete function with package name, necessary imports, and any required type definitions. Do not repeat docstring provided to you. Do not add additional explanations, or a main function.
# User
[Instruction]
Table 19: Prompt Template for Java Benchmark Runner in Code2Bench
@@
Table 20: Prompt Template for TypeScript Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write TypeScript code.
## Input The user provides you with an function signature and docstring, you should generate a TypeScript function based on them.
## Output
“‘ts
The generated TypeScript code.
## Note
- Only output self-contained TypeScript code without docstring and any additional information.
# User
[Instruction]
Table 21: Prompt Template for JavaScript Benchmark Runner in Code2Bench
# System
You are an expert in the field of coding, helping users write JavaScript code.
## Input
The user provides you with an function signature and docstring, you should generate a JavaScript function based on them.
## Output
“‘js
The generated JavaScript code.
## Note
- Only output JavaScript code without docstring and any additional information.
# User
[Instruction]
Appendix I Detailed Evaluation
I.1 Compute Resources
we detail the compute resources utilized for evaluating the Large Language Models on the Code2Bench-2505 benchmark. The evaluation was conducted on an infrastructure consisting of server-grade machines. Open-source models with fewer than 32B parameters (as listed in Table 2) were served using vLLM on a cluster equipped with NVIDIA GPUs. Specifically, these models were evaluated on machines featuring NVIDIA A100 80GB GPUs. The evaluation environment for these models was containerized to maintain isolation and consistency. For larger open-source models (>= 32B parameters) and all closed-source models, evaluation was performed by accessing their respective official APIs. The compute resources for these API-based evaluations are managed by the model providers and are not under our direct control or knowledge. Therefore, we cannot provide specific details on the underlying hardware, memory, or parallelization used by these providers. The Testcase Runner execution for each task (which involves loading test cases, running the generated code and ground truth, and performing differential testing) was primarily CPU-bound and ran on standard server CPUs, such as Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz, featuring 64 logical cores. These machines were equipped with 125 GiB of RAM and SSD storage for the benchmark data and test cases. The total compute time required for the comprehensive evaluation of all 16 models across the 1163 tasks in CODE2BENCH-2505 was substantial. While precise timing varies per model and task, we estimate the total GPU-hours consumed for the open-source model inferences to be approximately 200 GPU-hours. The total CPU-hours consumed for Testcase Runner execution across all models (including running the generated code and ground truth against approximately 500 tests per task per model) is estimated to be approximately 200 CPU-hours.
I.2 Detailed Performance and Failure Analysis
This appendix provides a detailed breakdown of the performance and failure type distribution for each evaluated model across all task types (Self-Contained tasks in Python, JS, TS, Java, Go, and Weakly Self-Contained tasks in Python). The results for each model and task type are presented as percentages of the total number of tasks in that category, broken down into successful completions (Success), completions resulting in execution errors (ExecutionError), and completions failing test cases (TestFailure). The sum of these three percentages for each task type and model is approximately 100%. For clarity, the data is split across multiple tables.
Table 22: Detailed Performance and Failure Type Distribution (Part 1). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
Doubao-1-5-pro-32k 36.27 6.74 56.99 27.98 28.50 43.52 25.91 10.36 63.73 35.23 5.70 59.07 34.72 5.18 60.10 53.02 9.09 37.89
Gemini-2.5-flash 39.90 7.77 52.33 27.46 28.50 44.04 28.50 16.06 55.44 34.20 7.77 58.03 35.23 7.77 56.99 53.02 8.08 38.90
Gemma-3-27b-it 34.20 5.70 60.10 21.24 30.57 48.19 22.80 15.54 61.66 28.50 10.88 60.62 29.53 10.88 59.59 45.96 8.08 45.96
Mistral-small-3.1 31.61 9.84 58.55 12.95 51.30 35.75 22.80 17.10 60.10 23.83 24.87 51.30 30.57 5.70 63.73 44.44 11.62 43.94
Table 23: Detailed Performance and Failure Type Distribution (Part 2). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
LLAMA-3.2-3B 18.13 19.69 62.18 3.11 77.20 19.69 6.74 33.16 60.10 7.77 34.20 58.03 9.33 37.82 52.85 23.23 23.74 53.03
LLAMA-4-scout 31.09 7.25 61.66 6.22 78.24 15.54 17.10 29.53 53.37 27.98 16.06 55.96 25.39 10.88 63.73 42.93 11.11 45.96
LLAMA-4-maverick 39.38 6.22 54.40 11.40 66.32 22.28 22.80 19.69 57.51 30.60 30.57 38.83 34.21 23.32 42.47 53.54 7.58 38.89
LLAMA-3.3-70B 37.31 7.25 55.44 15.54 59.07 25.39 24.35 18.65 56.99 27.46 15.03 57.51 27.46 9.33 63.21 45.45 8.08 46.46
Table 24: Detailed Performance and Failure Type Distribution (Part 3). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
Qwen3-235b-a22b 37.82 6.22 55.96 22.28 39.38 38.34 27.46 13.99 58.55 33.16 8.29 58.55 29.53 6.74 63.73 50.00 9.60 40.40
Qwen3-32b 35.23 7.25 57.51 10.36 48.70 40.93 24.87 14.51 60.62 30.05 7.77 62.18 27.98 7.77 64.25 45.45 11.62 42.93
Qwen2.5-Coder-14b 31.61 8.81 59.59 16.58 55.96 27.46 22.80 12.95 64.25 29.53 9.33 61.14 29.53 9.33 61.14 43.43 16.16 40.40
Qwen2.5-Coder-32b 36.27 8.29 55.44 17.10 59.07 23.83 24.35 12.95 62.69 32.12 6.74 61.14 34.20 10.88 54.92 44.44 14.65 40.91
Table 25: Detailed Performance and Failure Type Distribution (Part 4). S=Success, E=Exec. Error, F=Test Failure.
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
S E F S E F S E F S E F S E F S E F
Qwen3-8b 28.50 17.62 53.89 11.92 60.10 27.98 17.10 17.10 65.80 21.24 9.84 68.91 20.73 10.88 68.39 39.90 19.19 40.91
Qwen3-1.7b 14.51 22.28 63.21 1.55 87.56 10.88 9.33 46.63 44.04 9.33 21.76 68.91 10.36 20.21 69.43 20.20 31.82 47.98
Claude-3.7-sonnet 39.90 5.18 54.92 21.24 39.90 38.86 29.53 12.95 57.51 35.23 5.18 59.59 33.68 5.18 61.14 48.99 10.10 40.91
DeepSeek-V3 39.38 6.74 53.89 23.83 34.20 41.97 25.91 12.95 61.14 32.64 5.70 61.66 34.20 6.22 59.59 50.51 9.09 40.40
I.3 Detailed Performance by Difficulty Levels
This subsection provides the detailed average Pass@1 performance across evaluated models for SC tasks in different languages (Python, Go, TS, JS) and WSC tasks in Python, broken down by difficulty levels (Easy, Medium, Hard). The average Pass@1 is calculated as the sum of the accuracy values for all models within that difficulty level and language/task type, divided by the number of evaluated models (16 in total).
Table 26: Average Pass@1 Performance by Difficulty Level across Languages and Task Types (%).
Language/Task Type Easy (%) Medium (%) Hard (%)
SC Python 41.94 28.42 28.03
SC Go 19.75 16.50 6.40
SC TS 28.99 27.83 21.05
SC JS 30.43 28.34 20.72
WSC Python 65.75 36.11 15.12
The data presented here correspond to the aggregated results used for calculating the average Pass@1 per difficulty level shown in Table 4 in the main paper. The total number of task instances and correct completions contributing to these percentages are also shown in the raw data (provided alongside the benchmark release). This detailed breakdown allows for a finer-grained analysis of how model performance varies with the inherent complexity of the tasks within each language and task type category.
I.4 Detailed Pass Rate $\geq$ 98% Distribution
Table 27: Percentage of Problems with Test Case Pass Rate $\geq$ 98% for Evaluated Models across Task Types (%).
Model Name SC Python SC Go SC Java SC TS SC JS WSC Python
Claude-3.7-sonnet 13.47 14.69 15.58 14.18 11.40 10.10
Doubao-1-5-pro-32k-250115 12.44 13.57 14.94 9.84 11.40 9.09
Gemini-2.5-flash-preview 11.92 8.52 9.33 10.03 10.88 8.08
DeepSeek-V3 11.92 9.51 16.47 10.88 12.44 9.09
Llama-4-maverick 5.70 8.38 13.75 6.74 7.18 8.58
Qwen3-235b-a22b 12.95 10.79 10.39 10.36 12.44 9.60
Llama-4-scout 13.99 11.52 12.87 7.77 10.09 11.11
Llama-3.3-70b-Instruct 11.40 7.15 12.33 8.81 9.84 8.08
Qwen3-32b 9.84 8.56 12.64 9.33 11.92 11.62
Qwen2.5-Coder-32b-Instruct 8.81 7.65 13.79 9.84 8.81 14.65
Gemma-3-27b-it 8.81 7.72 9.95 10.88 9.33 8.08
Mistral-small-3.1-24b-Instruct 5.70 7.91 10.59 8.29 8.81 11.62
Qwen2.5-Coder-14b-Instruct 10.36 8.73 11.49 10.02 11.40 16.16
Qwen3-8b 9.84 9.91 11.34 10.88 10.88 19.19
Llama-3.2-3b-Instruct 9.54 4.17 3.59 6.86 4.93 23.74
Qwen3-1.7b 8.29 5.51 6.19 6.84 7.25 31.82
This subsection provides the detailed percentage of problems where evaluated models achieved a test case pass rate of 98% or higher for all evaluated models across all task types (SC Python, SC Go, SC Java, SC TS, SC JS, and WSC Python). This metric serves as an indicator of how frequently models produce code that is nearly correct but fails on a small number of subtle test cases generated by Property-Based Testing.
As demonstrated in Table 27, the percentage of problems where models achieved a pass rate of 98% or higher varies significantly across models and task types, generally ranging from less than 1% to over 30%. These results underscore the discriminative power of the rigorous PBT-generated test suites in identifying subtle errors that models frequently make, even when the generated code appears mostly correct.
Appendix J Limitations
Despite its strengths in generating dynamic, rigorously tested, and realistic tasks focusing on functional correctness, Code2Bench-2505, like many benchmarks, has limitations in its evaluation scope. Our primary focus is on assessing the functional correctness of generated code, measured through Pass@1 against comprehensive PBT-generated test suites. While functional correctness is paramount, real-world software development necessitates evaluating other crucial aspects of code quality, such as efficiency, readability, style, security, robustness to invalid inputs, and the ability to generate accompanying documentation or tests. Code2Bench-2505 currently does not directly evaluate these important dimensions.
Furthermore, the current iteration of Code2Bench-2505 primarily focuses on code generation tasks, where models are required to generate a complete function implementation based on a natural language instruction and function signature. However, the underlying structure of the benchmark, including the availability of ground truth implementations and the rigorous, diverse test cases generated via PBT, offers significant potential for evaluating LLM capabilities beyond simple generation. By leveraging the ground truth and PBT-generated test suites, the framework could be extended to support other task types crucial for software development workflows, such as code completion (filling in missing parts of code), code editing/repair (modifying existing code to meet new requirements or fix bugs), and assessing code reasoning abilities through execution prediction or debugging tasks. Expanding to these diverse task types would provide a more comprehensive evaluation of LLMs’ understanding and manipulation of code, moving beyond pure synthesis.

### v2

@@ I.1 Compute Resources
I.2 Evaluation Instruction
Table 15: Prompt Template for SC-Python Benchmark Runner in Code2Bench
@@ I.2 Evaluation Instruction
Table 16: Prompt Template for WSC-Python Benchmark Runner in Code2Bench
@@
Table 17: Prompt Template for SC-Java Benchmark Runner in Code2Bench
@@
Appendix J Case Study: Uncovering the Illusion of Correctness
Our diagnostic approach, combining a scaled source of real-world problems with the scaled rigor of Property-Based Testing, allows us to move beyond simple pass/fail metrics and uncover nuanced failure modes. This case study on a Weakly Self-Contained (WSC) task from CODE2BENCH-2509 illustrates how our framework reveals the critical gap between functional plausibility and engineering robustness.
J.1 The Task: A Numerically Sensitive Problem
WSC Task #81 requires the implementation of a _first_divided_difference function, a common operation in numerical analysis. The ground-truth implementation, sourced from a mature scientific Python library, is a highly efficient and numerically stable vectorized solution using NumPy:
1# Ground Truth (Vectorized, Numerically-Stable)
2def _first_divided_difference(d, fct, fctder, atol=1e-12, rtol=1e-12):
3 dif = np.repeat(d[None, :], len(d), axis=0)
4 close_ = np.isclose(dif, dif.T, atol=atol, rtol=rtol)
5 dif[close_] = fctder(dif[close_])
6 dif[~close_] = (fct(dif[~close_]) - fct(dif.T[~close_])) / \
7 (dif[~close_] - dif.T[~close_])
8 return dif
J.2 The “Near-Perfect” but Flawed LLM Solution
Remarkably, nearly all 10 evaluated models failed this task in the exact same way. They did not produce syntax errors or obvious logical flaws. Instead, they generated a functionally plausible solution that mimics a textbook implementation using scalar Python loops:
1# Typical LLM-Generated Solution (Scalar, Naive)
2def _first_divided_difference(d, fct, fctder, atol=1e-12, rtol=1e-12):
3 n = len(d)
4 fdd = np.zeros((n, n))
5 for i in range(n):
6 for j in range(n):
7 if np.isclose(d[i], d[j], atol=atol, rtol=rtol):
8 fdd[i, j] = fctder(d[i])
9 else:
10 fdd[i, j] = (fct(d[i]) - fct(d[j])) / (d[i] - d[j])
11 return fdd
The diagnostic power of our benchmark is revealed in how this seemingly correct solution failed. The code generated by DeepSeek-V3 for this task passed an astonishing 98.8% of our PBT-generated test cases (494 out of 500). It only failed on a few specific, numerically challenging inputs where the different order of floating-point operations between the vectorized and scalar approaches led to minute rounding errors. These tiny discrepancies, while functionally insignificant in many contexts, were caught by our strict-tolerance deep comparison function. For example, one failing test case reported:
> Mismatch found. Expected: ...219e-08, Actual: ...224e-08
J.3 Actionable Insights from a “Near-Miss” Failure
This single “near-miss” failure pattern, consistent across the entire model spectrum, provides several highly actionable insights that would be invisible to conventional benchmarks:
  • For LLM Developers: This reveals that models learn to be academically correct, but not industrially robust. They successfully reproduce textbook patterns but lack essential engineering knowledge regarding idiomatic code (vectorization), performance optimization, and numerical stability. To close this gap, training data should be augmented to explicitly reward these non-functional properties. Our benchmark, with its real-world ground truths and precision-sensitive tests, provides ideal data for such targeted fine-tuning.
  • For Benchmark Designers: This case powerfully validates our deep testing approach and exposes the limitations of shallow test suites. A typical benchmark, likely using only a few simple integer-based test cases, would have falsely labeled this numerically unstable solution as a success. Only through the exhaustive, edge-case-driven nature of our PBT methodology is the critical difference between a “toy” solution and a robust one revealed—penalizing brittle, “good-enough” outputs and rewarding true engineering rigor.
This example epitomizes the diagnostic philosophy of CODE2BENCH: to not just reveal what fails, but to provide deep insights into why it fails and how future models can be improved.
Appendix K Scalability
K.1 Advanced Extensibility: Hierarchical Dependency Resolution
The dependency classification into SC and WSC, as described in the main paper, provides a robust foundation for benchmark curation. However, a significant portion of real-world code involves functions that call other project-internal functions. While these are classified as Project-Dependent (PD) and typically discarded, a substantial subset of them are, in fact, hierarchically testable. This observation opens a powerful new avenue for Scaling the Source even further.
K.1.1 The Concept of Layered Self-Contained (LSC) Tasks
We define a Layered Self-Contained (LSC) function as a function that is not strictly SC or WSC itself, but whose entire set of project-internal dependencies recursively resolves to a set of functions that are all either SC or WSC.
Consider a function $f_{A}$ that calls another internal function $f_{B}$.
  • If $f_{B}$ is Self-Contained (SC), then the functional behavior of $f_{A}$ is fully determined by its own logic and the well-defined, dependency-free logic of $f_{B}$.
  • Similarly, if $f_{B}$ is Weakly Self-Contained (WSC), the behavior of $f_{A}$ is determined by its logic and the behavior of $f_{B}$, which itself is only dependent on a set of allowed public libraries.
In both cases, the complete functional behavior of the top-level function $f_{A}$ can be fully specified and is not reliant on any un-testable, opaque, or proprietary internal state. Therefore, it is a suitable candidate for a rigorous, standalone benchmark task.
K.1.2 Methodology for LSC Task Generation and Verification
Our CODE2BENCH framework can be extended to identify and generate these LSC tasks through a recursive dependency analysis powered by our Scope Graph:
  1. 1.
    Recursive Dependency Resolution: When a function $f_{A}$ is initially classified as PD due to a call to an internal function $f_{B}$, our framework does not immediately discard it. Instead, it recursively runs the dependency analysis on $f_{B}$. This process continues until all dependencies are either resolved to primitives, allowed libraries, or the dependency chain terminates.
  2. 2.
    Hierarchical Test Oracle Construction: To create a test oracle for an LSC function like $f_{A}$, we provide not only its own source code but also the source code of its entire dependency tree of SC/WSC functions (e.g., $f_{B}$, and any functions $f_{B}$ calls). This complete, self-contained bundle of functions serves as the ground-truth implementation.
  3. 3.
    PBT-based Verification: Property-Based Testing is then applied to the top-level function $f_{A}$. The PBT engine generates inputs for $f_{A}$, and the complete, bundled ground-truth implementation is executed to generate the expected outputs. The 100% branch coverage quality gate is applied to this entire bundle, ensuring that the tests thoroughly exercise not only the logic of $f_{A}$ but also the interactions with its internal dependencies.
This extension to handle LSC tasks dramatically increases the pool of high-quality, testable functions that can be extracted from real-world repositories. It allows our framework to capture more complex, multi-function interactions that are representative of real-world software design, while still maintaining the rigorous, deterministic verifiability that is the hallmark of our approach. This represents a significant future direction for scaling the realism and complexity of the CODE2BENCH suite.
K.2 Generating Syntax-Aware Code Completion Tasks
A key design principle of the CODE2BENCH framework is its extensibility beyond single-function generation. The true assets curated by our pipeline are the large collection of high-quality, real-world ground-truth functions, each paired with a comprehensive suite of high-coverage Property-Based Tests. This powerful combination of a "solution" ($f_{\text{gt}}$) and a rigorous "verification" mechanism (the PBT suite) provides a uniquely powerful foundation for generating a wide array of challenging and realistic software engineering tasks. In this section, we detail how our framework can be systematically extended to Code Completion.
Our framework can automatically generate high-quality, syntax-aware "fill-in-the-middle" code completion tasks. Inspired by prior work on structured code completion Gong et al. (2024), we can leverage our curated SC and WSC function pools to create distinct types of completion challenges:
  • Completion-SC (Algorithmic Logic Completion): For our Self-Contained (SC) tasks, which are rich in algorithmic logic, we can create completion benchmarks by masking entire logical blocks.
  • Completion-WSC (API Call Completion): For our Weakly Self-Contained (WSC) tasks, which are centered on library usage, we can create completion benchmarks by masking specific API calls.
K.3 Rigorous Verification via PBT
The most significant advantage of deriving completion tasks from CODE2BENCH is the automatic inheritance of our rigorous verification mechanism. Unlike many completion benchmarks that rely on simple syntactic checks (e.g., exact match or BLEU score), we can evaluate the functional correctness of the completed code.
K.4 Functional Correctness vs. Robustness Trade-off.
As highlighted by reviewers, our stringent 100% branch coverage gate effectively filters out complex defensive logic (e.g., unreachable error handling branches) that is difficult to trigger via random inputs. This represents a deliberate strategic choice: we prioritized establishing an unimpeachable “gold standard” for core algorithmic and logical correctness over the coverage of defensive programming constructs. While this ensures absolute verifiability, it temporarily de-emphasizes the evaluation of code robustness. However, our framework is designed to retrieve this filtered data. In future work, we plan to construct a dedicated “Robustness Benchmark” by tasking models to add comprehensive error handling (e.g., try-catch, input validation) to the verifiable “happy path” implementations curated in this work.
K.5 Beyond Function-Level Isolation.
The current iteration focuses on function-level tasks to ensure unit-testable rigor. We acknowledge that real-world software engineering involves complex, project-level dependencies. Crucially, our framework is methodologically ready for this expansion. Our Scope Graph analysis (§LABEL:sec:method) natively supports resolving cross-file dependencies and class hierarchies. By combining this with Stateful Property-Based Testing (Stateful PBT), which can generate sequences of API calls rather than static data, we envision scaling our rigorous verification pipeline to Project-Dependent (PD) tasks that assess multi-function interactions and state management.
K.6 Multi-Dimensional Evaluation.
Our primary metric is Pass@1 based on functional correctness. While fundamental, this does not directly measure other code quality attributes such as efficiency, security, or readability. However, our massive suite of ground-truth functions and generated test cases serves as a versatile asset. Future iterations can leverage these assets to benchmark execution time (efficiency), check against secure coding standards (security), or serve as references for style compliance (readability), providing a more holistic assessment of LLM coding capabilities.
Appendix L Limitations
Despite its strengths in generating dynamic, rigorously tested, and realistic tasks focusing on functional correctness, Code2Bench-2509, like many benchmarks, has limitations in its evaluation scope. Our primary focus is on assessing the functional correctness of generated code, measured through Pass@1 against comprehensive PBT-generated test suites. While functional correctness is paramount, real-world software development necessitates evaluating other crucial aspects of code quality, such as efficiency, readability, style, security, robustness to invalid inputs, and the ability to generate accompanying documentation or tests. Code2Bench-2509 currently does not directly evaluate these important dimensions.
Furthermore, the current iteration of Code2Bench-2509 primarily focuses on code generation tasks, where models are required to generate a complete function implementation based on a natural language instruction and function signature. However, the underlying structure of the benchmark, including the availability of ground truth implementations and the rigorous, diverse test cases generated via PBT, offers significant potential for evaluating LLM capabilities beyond simple generation. By leveraging the ground truth and PBT-generated test suites, the framework could be extended to support other task types crucial for software development workflows, such as code completion (filling in missing parts of code), code editing/repair (modifying existing code to meet new requirements or fix bugs), and assessing code reasoning abilities through execution prediction or debugging tasks. Expanding to these diverse task types would provide a more comprehensive evaluation of LLMs’ understanding and manipulation of code, moving beyond pure synthesis.
@@
Our evaluation focuses on a diverse suite of state-of-the-art, instruction-following code generation models. We acknowledge that our study does not include models specifically designed or fine-tuned for multi-step, competitive-programming-style reasoning (e.g., models employing complex search algorithms or Chain-of-Thought prompting for code). This exclusion was a deliberate choice based on two primary considerations. First, our preliminary explorations indicated that the verbose, multi-step reasoning outputs of such models often exceeded practical token limits for our large-scale, automated evaluation harness, presenting significant computational and financial costs. Second, and more critically, the primary goal of CODE2BENCH is to evaluate a model’s ability to generate direct, production-style code from real-world specifications, a task for which current instruction-following models are the most direct fit. While evaluating deep reasoning capabilities is an important research direction, it represents a different evaluation paradigm that is beyond the scope of our current study.